A sample of nitrogen gas has a volume of 3.76 L at 27°C. What volume will it occupy at 270. K if the pressure and number of mol are constant? Enter your answer in the provided box.

1
Expert's answer
2021-10-05T23:58:37-0400

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{3.76}{(27+273)}=\frac{V_2}{270}"


"\frac{3.76}{(300)}=\frac{V_2}{270}"


"V_2=\frac{3.76×270}{(300)}=3.384litres"